OpenStudy (anonymous):

please help me solve: (x-4)-5(x-4)^(1/2)=6 thanks!

myininaya (myininaya):

i remember doing this one

myininaya (myininaya):

let me see if i can find it lol

OpenStudy (anonymous):

me too. your turn

OpenStudy (anonymous):

think of it as a quadatic equation in \[(x-4)^{\frac{1}{2}}\]

OpenStudy (anonymous):

\[z^2-5z-6=0\] \[(z-3)(z-2)=0\] \[z=3,z=2\] \[\sqrt{x-4}=3\] or \[\sqrt{x-4}=2\] etc

OpenStudy (anonymous):

the 'etc' being \[\sqrt{x-4}=3\] \[x-4=9\] \[x=13\] or \[\sqrt{x-4}=2\] \[x-4=4\] \[x=8\]

OpenStudy (anonymous):

so my answers are wrong???

myininaya (myininaya):

you factored wrong

OpenStudy (anonymous):

lololololololol

myininaya (myininaya):

(x-3)(x-2) does not equal x^2-5x-6 (x-6)(x+1)=x^2-5x-6

OpenStudy (anonymous):

\[(z-6)(z-1)=0\] \[z=6,z=-1\] \[\sqrt{x-4}=6\] \[x-4=36\] \[x=40\] whew

myininaya (myininaya):

lol

OpenStudy (anonymous):

time to cook